### v2

# Verbosity-Aware Rationale Reduction: Effective Reduction of Redundant Rationale via Principled Criteria

Abstract
Large Language Models (LLMs) rely on generating extensive intermediate reasoning units (e.g., tokens, sentences) to enhance final answer quality across a wide range of complex tasks. While generating multiple reasoning paths or iteratively refining rationales proves effective for improving performance, these approaches inevitably result in significantly higher inference costs. In this work, we propose a novel sentence-level rationale reduction training framework that leverages likelihood-based criteria, verbosity, to identify and remove redundant reasoning sentences. Unlike previous approaches that utilize token-level reduction, our sentence-level reduction framework maintains model performance while reducing generation length. This preserves the original reasoning abilities of LLMs and achieves an average 17.15% reduction in generation costs across various models and tasks.
Verbosity-Aware Rationale Reduction:
Effective Reduction of Redundant Rationale via Principled Criteria
Joonwon Jang1  Jaehee Kim 2  Wonbin Kweon3  Hwanjo Yu1,4,∗ POSTECH1,3,4, Seoul National University2 {kaoara, skwb4453, hwanjoyu}@postech.ac.kr 1,3,4 jaehee_kim@snu.ac.kr 2
1 Introduction
Recent advances in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities comparable to human cognitive abilities (Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). These models can solve complex reasoning tasks by explicitly generating extended reasoning paths. The generation of such paths involves producing explicit reasoning units (e.g., tokens, steps) (Yu et al., 2024b), which additionally enhances model performance through iterative prompting (Wang et al., 2023; Yao et al., 2023). This iterative generation of explicit reasoning paths allows the model to refine and expand its thought processes, incorporating strategic planning and ongoing cognitive generation (Xi et al., 2023; Yang et al., 2024).
Although the extensive generation of explicit reasoning units leads to better peformance, it inherently leads to higher inference costs and increased latency (Yu et al., 2024b; Wang et al., 2024). Moreover, fine-tuning LLMs using a complete reasoning path does not consistently guarantee stronger performance (Yu et al., 2024b; Deng et al., 2024b; Liu et al., 2024), suggesting the need for methods that can maintain reasoning performance while reducing the generation of reasoning units. Despite this apparent requirement, it remains underexplored how to maintain LLM reasoning capabilities while reducing intermediate reasoning paths across diverse tasks.
Previous methods (Yu et al., 2024b) have drawn inspiration from human cognitive processes to address the aforementioned limitations. Some studies have proposed training pipelines that utilize augmented datasets, iteratively generated by foundation LLMs, to fine-tune subsequent LLMs (Yu et al., 2024b; Liu et al., 2024). However, they are inherently vulnerable as they significantly depend on the generative capabilities of LLMs. In response, other works have directly trained LLMs without augmenting datasets that facilitate implicit reasoning before answer generation. Deng et al. (2023) introduced a knowledge distillation method to distill explicit reasoning into implicit reasoning through token-level hidden states. Deng et al. (2024b) selected tokens as the reasoning unit for reduction and proposed a heuristic method to internalize explicit intermediate rationale tokens. They progressively eliminate tokens from the beginning of the full rationale during the Chain of Thought (CoT) fine-tuning. However, these methods present a fundamental limitation as they lack sufficient justification for selecting tokens over more linguistically natural units (e.g., sentences) for reduction (Table 1), and they fail to provide principled criteria for the removal process. Furthermore, their evaluation has been predominantly conducted on synthetic arithmetic reasoning tasks, which presents limitations for real-world applications.
To address these limitations, we propose a novel training method for maintaining LLM reasoning performance while systematically eliminating redundant reasoning units within the reasoning process. Our method leverages sentences as fundamental reduction units, establishing more linguistically coherent boundaries compared to token-level approaches. We first empirically demonstrate that rationale sentences in early reasoning stages can introduce redundancy in LLM’s answer generation process. Drawing inspiration from Dong et al. (2023), we introduce the concept of ‘verbosity’ and develop a CoT fine-tuning method that removes unnecessary rationale sentences, thereby reducing rationale generation while preserving reasoning capabilities.
The primary contributions of this work are as follows:
  1. 1.
    We propose a novel method for sentence-level reasoning path reduction in LLMs.
  2. 2.
    We demonstrate that principled criteria can effectively identify redundancies and reduce intermediate token generation efficiently.
  3. 3.
    We validate our method’s effectiveness and robustness through extensive experiments across various LLMs and ablation studies.
2 Related Works
2.1 Performance-Cost Tradeoffs in Reasoning Path Generation
Recent research has demonstrated the importance of generating iterative and refined reasoning paths to improve the reasoning capabilities of models, albeit at increased computational costs (Wang et al., 2023; Yao et al., 2023; Radha et al., 2024; Wang et al., 2024; Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024). Self-Consistency (Wang et al., 2023) improves reasoning accuracy by generating more than 20 distinct reasoning paths and selecting the most consistent answer using ensemble methods, resulting in substantial computational overhead. Similarly, Tree of Thoughts (ToT) (Yao et al., 2023) expands the exploration of reasoning paths using a tree-structured method, while Strategic Chain of Thought (SCoT) (Wang et al., 2024) utilizes an Inner Dialogue Agent to adaptively explore the diverse reasoning path.
Concurrently, Self-Refine (Madaan et al., 2024) utilizes an iterative framework that incorporates feedback to refine responses and enhance reasoning accuracy, necessitating multiple forward passes through the model. Additionally, the Reflexion-based framework (Shinn et al., 2024; Kumar et al., 2024) guides models to iteratively generate relevant reasoning paths through reflection for accurate answer generation. Given the necessity of iterative generation and refinement of reasoning paths to achieve optimal performance, they inherently increase inference costs and latency. Therefore, it is essential to investigate methods for efficiently generating these paths.
2.2 Reasoning Path Reduction
To address the computational costs associated with extensive reasoning paths generation, some lines of work (Yu et al., 2024b; Liu et al., 2024) have focused on generating augmented datasets with varying rationale lengths to shorten the generation of reasoning paths. Yu et al. (2024b) guide model to generate multiple reasoning paths using Self-Consistency to construct datasets, subsequently fine-tuning the model to generate direct answers without intermediate reasoning steps. Liu et al. (2024) developed a heuristic approach to merge reasoning steps and iteratively trained the model to produce shorter reasoning paths, which are subsequently incorporated into the progressive training phase. While these approaches demonstrate empirical effectiveness, they present two fundamental limitations: (1) their heavy dependence on LLM generation capabilities introduces inherent instability, and (2) their objective of reducing reasoning path necessitates the paradoxical creation of datasets requiring extensive reasoning path generation.
To address these weaknesses, another line of work (Deng et al., 2023, 2024b) has focused on directly training LLM without augmented datasets. Implicit-CoT (Deng et al., 2023) utilizes a multi-model framework where an emulator model is trained to predict the teacher’s token-level hidden states, and a student model uses these predicted states to generate answers. ICoT-SI (Deng et al., 2024b) identifies tokens as reduction units, proposing a method to internalize explicit intermediate rationale tokens by progressively removing them from the beginning of the reasoning path within the CoT fine-tuning process. However, these methods have shown limited generalization across diverse datasets as they have only been validated on simple arithmetic reasoning tasks, such as multiplication problems. This limitation raises concerns about their applicability to real-world scenarios where rationales are expressed in natural language. Moreover, they do not justify their criteria for token-level removal nor consider whether more linguistically natural units (e.g., sentences) might serve as more effective reduction units. Specifically, the token-level removal approach may inadvertently eliminate critical information necessary for answer generation or distort the semantic information of the sentence, while requiring repeated removal of redundant reasoning steps. Motivated by the computational costs of the extensive reasoning path generation and the limitations of existing reduction approaches, we investigate the redundancy of various sentence positions for potential removal and propose a novel method with principled criteria that can efficiently reduce reasoning paths while maintaining their effectiveness.
3 Early Stage Rationales are Redundant
Figure 1: NLL differences across varying sizes of $\{r_{i}\}$. The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that $\{r_{i}\}$ is sampled from the front, random, and back indices of the full index set, respectively.
3.1 Quanitifying the redundancy
Before delving into the method, we first investigate which positions within the rationale sentences should be selected for reduction. When the likelihood of the answer remains unchanged after a sentence is removed from the full rationales, it indicates that the sentence may be redundant for the reasoning process. To quantify the redundancy of a sentence, we compute the negative log-likelihood (NLL) for answer $y$ after the sentence reduction as follows:
$\begin{split}\text{NLL}&=-\log p_{\theta}(y|R^{\prime},x),\\ \text{where}\ R^{\prime}&=R\setminus\{r_{i}\}_{i\in S},\ S\subseteq I.\end{split}$ (1)
Let $R$ denote the complete set of rationale sentences, and $I$ represent the full index set of these sentences. The subset of indices corresponding to sentences selected for reduction is denoted by $S$, and $R^{\prime}$ represents the remaining rationale sentences after their removal. $\{r_{i}\}_{i\in S}$ denotes the sentences selected for reduction. For simplicity, we use $\{r_{i}\}$ without the subset index notation throughout the rest of the paper.
3.2 Redundancy of leading sentences
We conduct a pilot study to empirically demonstrate the redundancy of the leading sentences within the rationales. By varying the size of $\{r_{i}\}$ from 1 to 4, we calculate the NLL of Mistral 7B (Jiang et al., 2023) across different reasoning datasets. To explore more general patterns of sentence removal, we employ a stochastic strategy in our experimental setup. We compare three configurations for sampling $\{r_{i}\}$: front, random, and back. In the front configuration, initial sentences are assigned a higher probability when forming $R^{\prime}$, whereas random and back configurations use uniform and progressively increasing probabilities, respectively [^1] For generating $R^{\prime}$, we assign probabilities $p_{k}=\frac{N-k+1}{\sum_{i=1}^{N}i}$ (front), $\frac{1}{N}$ (random), and $\frac{k}{\sum_{i=1}^{N}i}$ (back) where k=1,…,N denotes sentence position.. Additionally, we computed the NLL for full rationale sentences (i.e., $-\log p_{\theta}(y|R,x)$) as a baseline to evaluate the impact of reduction. As depicted in Figure 1, the NLL difference between the full and reduced rationales shows a marginal likelihood difference in the front ($<$) configuration. In contrast, removing sentences randomly ($\diamond$) or from the back ($>$) results in higher NLL as the removed sentences increase, highlighting the importance of the selection of removal position in the reasoning and answer prediction process (for additional data, see Appendix A).
4 Verbosity-Aware Rationale Reduction
Figure 2: Overview of our VARR/VARR+ framework. Initially, we select a candidate sentence from the beginning of the rationale (Section 3). After removing the candidate sentence, we evaluate $\textit{verbosity}(y_{g})$ and $\textit{verbosity}(y_{w})$ using Equations (5) and (9). If the candidate sentence meets the verbosity evaluation criteria, it is excluded in subsequent training steps. The model then proceeds with training, where the redundant sentence is excluded from the rationale.
Based on the above observation, we propose the Verbosity-Aware Rationale Reduction (VARR) framework. In Section 4.1, we introduce the concept of ‘verbosity’ as a principled criterion for identifying redundant reasoning sentences. Section 4.2 describes the integration of verbosity in the reduction process during CoT training. Subsequently, Section 4.3 presents additional verbosity term by incorporating the wrong answer to improve robustness. Finally, Section 4.4 illustrates the overall VARR framework.
4.1 Verbosity
To quantify the redundancy of the sentences for removal, we first introduce a foundational concept ‘verbosity’. Given an input $x$, full rationale $R$, and a reduced rationale $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, we measure the verbosity of a sentence $r_{i}$ on $y$ by measuring the difference in Kullback–Leibler divergence (KL-divergence) as follows:
$\begin{split}\textit{verbosity}(y)&=D_{KL}(q(y|x)\parallel p_{\theta}(y|R,x))\\ &\hskip 15.00002pt-D_{KL}(q(y|x)\parallel p_{\theta}(y|R^{\prime},x)),\end{split}$ (2)
where $q(y|x)$ is the ground truth distribution. The verbosity$(y)$ highlights the informational contribution or redundancy of a rationale sentence $r_{i}$ on answer $y$. Since $q(y|x)$ is the form of the one-hot vector (i.e., Dirac delta function), the verbosity$(y)$ can be computed as the log-likelihood ratio between $R$ and $R^{\prime}$ as follows:
$\begin{split}\textit{verbosity}(y_{g})&=\left[H_{q}(p_{\theta}(y|R,x))-H(q(y|x))\right]\\ &\hskip 15.00002pt-\left[H_{q}(p_{\theta}(y|R^{\prime},x))-H(q(y|x))\right]\\ &=E_{q}[-\log P(y|R,x)]\\ &\hskip 15.00002pt+E_{q}[\log P(y|R^{\prime},x)]\\ &=\log\left(\frac{P(y_{g}|R^{\prime},x)}{P(y_{g}|R,x)}\right),\end{split}$ (3)
where $y_{g}$ is the ground truth answer (i.e., $q(y_{g}|x)=1$). $H_{q}(\cdot)$ and $H(\cdot)$ denote the cross-entropy and the entropy, respectively. Intuitvely, a higher value of $\textit{verbosity}(y_{g})$ implies that the likelihood of the model generating the ground truth answer increases after removing $r_{i}$, indicating that its removal is beneficial.
4.2 Verbosity Identification in CoT Training
Given an input sequence, the CoT training (Nye et al., 2021) aims to train LLMs to generate full rationale, followed by the ground truth answer with the following loss:
$-\log p_{\theta}(y_{g},R|x).$ (4)
In each training step $t$, each sentence $r_{i}$ within $R$ is assessed with the following the equation:
$\textit{verbosity}(y_{g})\geq 0.$ (5)
Here, we sequentially select $r_{i}$ starting from the first sentence and construct $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, based on our analysis in Section 3, which indicates that early-stage rationales are more likely to be redundant. A non-positive verbosity score indicates that removing $r_{i}$ from $R$ impairs the model’s performance, thereby identifying the sentence as essential and retaining it in the rationale.
4.3 Contrasting with Wrong Answer
Inspired by the miscalibrated log-likelihoods between accepted and rejected responses during standalone Supervised Fine-Tuning (SFT) in alignment learning Rafailov et al. (2024); Azar et al. (2024); Hong et al. (2024), we also assess whether the reduced rationales $R^{\prime}$ result in inaccuracy answer generation by introducing a wrong answer $y_{w}$. Instead of employing the correct answer distribution $q(y|x)$ in Equation (2), we initiate our formula with the wrong answer distribution $q^{\prime}$ (i.e., $1-q(y|x)$, normalized sum to 1) Through algebraic manipulation, we can obtain:
$\begin{split}&D_{KL}(q^{\prime}(y|x)\parallel p_{\theta}(y|R,x))\\ &\hskip 10.00002pt-D_{KL}(q^{\prime}(y|x)\parallel p_{\theta}(y|R^{\prime},x))\\ &=\left[H_{w}(p_{\theta}(y|R,x))-H(q^{\prime}(y|x))\right]\\ &\hskip 10.00002pt-\left[H_{w}(p_{\theta}(y|R^{\prime},x))-H(q^{\prime}(y|x))\right],\end{split}$ (6)
where $H_{w}(\cdot)$ denotes the cross-entropy calculated with the wrong answer distribution $q^{\prime}$. Due to the impracticality of computing the expectation over the entire space of $V-1$ wrong answers (where $V$ is the vocabulary size), we sample $K$ incorrect answers for the following estimations:
$\begin{split}&E_{w}[-\log p_{\theta}(y|R,x)]+E_{w}[\log p_{\theta}(y|R^{\prime},x)]\\ &={E}_{w}\Big{[}\log\frac{p_{\theta}(y|R^{\prime},x)}{p_{\theta}(y|R,x)}\Big{]}\\ &\approx\frac{1}{K}\sum_{k=1}^{K}\log\frac{p_{\theta}(y_{w}^{(k)}|R^{\prime},x)}{p_{\theta}(y_{w}^{(k)}|R,x)},\end{split}$ (7)
where $\{y_{w}^{(k)}\}_{k\in\left[K\right]}$ is sampled from the in-batch negatives depending on the dataset. Consequently, the $\textit{verbosity}(y_{w})$ is computed as:
$\textit{verbosity}(y_{w})=\frac{1}{K}\sum_{k=1}^{K}\log\frac{p_{\theta}(y_{w}^{(k)}|R^{\prime},x)}{p_{\theta}(y_{w}^{(k)}|R,x)}.$ (8)
Since computational constraints necessitate sampling incorrect answers to calculate $\textit{verbosity}(y_{w})$, we contrast $\textit{verbosity}(y_{w})$ with $\textit{verbosity}(y_{g})$ rather than solely using $\textit{verbosity}(y_{w})$. This ensures that the likelihood of incorrect answers remains lower than that of ground truth answers, even after the removal of $r_{i}$ as follows:
$\textit{verbosity}(y_{w})-\textit{verbosity}(y_{g})\leq 0.$ (9)
A smaller $\textit{verbosity}(y_{w})$ value compared to $\textit{verbosity}(y_{g})$ indicates that $R^{\prime}$ increases the likelihood of generating incorrect answers over the ground truth answer after removing $r_{i}$, making it beneficial to remove.
4.4 CoT Training with Rationale Reduction
In our framework, the model is trained using Equation (4) for predefined warm-up steps to inject its reasoning capabilities. After the warm-up steps, at each training step $t$, each sentence $r_{i}$, starting from the first index sentence in $R$, is evaluated against either Equation (5) alone (denoted as VARR) or both Equations (5) and (9) (denoted as VARR+). Sentences satisfying these respective criteria are selected for removal and are not included in subsequent training steps. The maximum removable number of sentences at each training step $t$ is determined based on a linear schedule, adopting Deng et al. (2024b)’s setting as follows:
$r(t)=\lfloor N_{t}\cdot(t/T)\rfloor,$ (10)
where $T$ represents the total number of training steps, $N_{t}$ is the total number of rationale sentences at step $t$, and $r(t)$ indicates the maximum number of sentences that can be removed at that step.
5 Experiments
5.1 Training Configuration
Figure 3: Performance for 5 different reasoning tasks. Acc refers to Accuracy while Tokens denotes the Generated Tokens. Bar plots illustrate the accuracy of each method on datasets, while line plots show the average generated tokens by each method across the datasets. The y-axis represents the range of each metric, and the x-axis lists the datasets evaluated.
Datasets
We conducted experiments across two categories to provide a comprehensive evaluation of VARR’s versatility and effectiveness, in contrast to prior research that predominantly focuses on simple arithmetic tasks like multi-digit multiplication (Deng et al., 2023, 2024b). Initially, we evaluate with arithmetic reasoning tasks, including datasets like GSM8K (G8K; Cobbe et al. 2021a) and MathQA (MQA; Amini et al. 2019a). We also examine the performance of our method on commonsense reasoning tasks, employing datasets including CommonsenseQA (CQA; Talmor et al. 2019), TrivaiQA (TQA; Joshi et al. 2017), and StrategyQA (SQA; Geva et al. 2021).
Models
We trained Mistral 7B (Jiang et al., 2023) as our base model for comparisons. We also trained a series of Llama3.2 models (AI@Meta, 2024) scaling from 1B to 3B to demonstrate our method’s generalization capabilities.
Implementation Details
The warm-up stage is set to 0.1 of the total training steps, and its impact is analyzed in Section 5.5.4. In addition, an optimizer is reinitialized at the beginning of each epoch to stabilize the model training inspired by Deng et al. (2024b)’s setting, with its effects described in Section 5.5.3. All baselines and our proposed method are trained for 5 epochs to ensure fairness in comparison. Detailed implementation of VARR/VARR+ can be found in the Appendix B.
5.2 Baselines
We compare our method to the following baselines: Explicit-CoT (Nye et al., 2021), where the model is finetuned with explicit chain-of-thought reasoning, and ICoT-SI (Deng et al., 2024b), where the model is fine-tuned using a linear token elimination schedule. It is important to note that after removing intermediate reasoning paths, the model’s performance should be competitive with Explicit-CoT. Consequently, ICoT-SI serves as our primary baseline. We exclude Implicit-CoT Deng et al. (2023) from our baselines due to its significant memory requirements, requiring the training of three models on a single GPU. Furthermore, according to Deng et al. (2024b), this method underperformed compared to ICoT-SI. All baselines were trained on a single A100-80GB GPU, and detailed training configurations for each method are provided in Appendix C.
5.3 Evaluation
We employ two evaluation metrics: First, we evaluate the accuracy of each method in generating the final answer for the respective tasks. Second, we count the generated tokens to evaluate reasoning efficiency while maintaining performance.
5.4 Main results
In Figure 3, we present the results for each reasoning task across different models. Our method demonstrates consistent performance improvements or maintains comparable results across most benchmarks, while also achieving more efficient reasoning path generation. Notably, we observed significant improvements in average accuracy with increments of 2.59%, 6.27%, and 14.27% for Mistral 7B, Llama3.2 1B, and Llama3.2 3B (VARR+), respectively, compared to Explicit CoT, while efficiency improved by 19.35%, 17.24%, and 14.88% for these models. These findings indicate that some reasoning sentences are redundant and can even disturb answer generation, highlighting that effective reasoning capabilities can be learned through selective utilization of essential reasoning units.
For ICoT-SI, the model exhibited performance degradation after token removal training. Specifically, ICoT-SI shows an average performance degradation of 18.93% compared to Explicit CoT. Furthermore, while the Mistral 7B model achieved a 50.26% accuracy on the GSM8K using a large augmented dataset with succinctly generated reasoning paths (trained with the 385K-size dataset used in the original paper (Deng et al., 2024b)), its performance dropped significantly to 28.27% when trained on the standard GSM8K dataset (trained with the 7K-size dataset). This indicates that while ICoT-SI may excessively reduce the length of rationales in reasoning tasks, identifying the appropriate reasoning unit and applying principled criteria are crucial for maintaining performance in real-world application problems. Qualitative results across baselines are presented in Appendix F, and a detailed analysis of the reasoning unit will be discussed in Section 5.5.1.
Moreover, introducing a wrong answer in VARR led to performance improvements in most datasets. While solely using the verbosity($y_{g}$) enhanced efficient generation, it is notable that verbosity($y_{w}$) constraints reduced rationales to increase the likelihood of generating incorrect answers, which contributes to the robustness and stability of the training process.
5.5 Ablation Studies
In this section, we conduct ablation studies to empirically validate our method. All experiments are implemented using Mistral 7B due to its higher base capacity compared to other models.
5.5.1 Identifying Appropriate Units for Removal
G8K MQA CQA TQA SQA
Exp-CoT 55.26 55.84 84.33 82.94 74.70
(122.54) (105.02) (42.84) (63.46) (46.47)
ICoT-SI 28.27 35.84 67.82 77.09 61.33
(104.41) (113.55) (6.0) (22.03) (22.86)
VARR-Tok 47.53 46.79 82.60 67.14 71.22
(94.16) (90.39) (25.07) (38.10) (31.92)
VARR-Sent 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 1: Analysis across various reasoning reduction units and the application of principled criteria. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
In this section, we experimentally demonstrate why the sentences are suitable units for reduction reasoning units. Unlike ICoT-SI which lacks specific criteria for token removal, we apply VARR+ at the token level (denoted as VARR-Tok) to assess whether tokens can effectively function as reduction units in reasoning processes. As depicted in Table 1, applying principled criteria at the token level results in an average performance gain of 26.95%, indicating that the principled criteria contribute to robust performance. Moreover, extending the reasoning units from tokens to sentences (denoted as VARR-Sent) leads to a further performance increase of 21.83% over VARR-Tok, highlighting the importance of selecting the appropriate reasoning unit for effective reasoning path reduction. These findings demonstrate the importance of selecting appropriate reasoning units, with sentences serving as linguistically natural and effective boundaries for the removal process.
5.5.2 Analyzing the Impact of Sentence Position on Removal Efficacy
G8K MQA CQA TQA SQA
Exp-CoT 55.26 55.84 84.33 82.94 74.70
(122.54) (105.02) (42.84) (63.46) (46.47)
No Rule 25.85 35.06 74.78 74.14 72.38
(63.56) (59.92) (23.56) (6.84) (23.38)
Random 52.31 55.34 83.47 79.62 73.83
(115.59) (99.13) (49.19) (44.98) (50.51)
Back 48.52 49.71 85.21 70.0 74.41
(103.25) (92.64) (19.86) (36.22) (18.17)
Front 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 2: Performance across 5 different reasoning tasks, evaluated with different sentence position removal. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
In Section 3, we demonstrated that gradually removing sentences from random and back positions can degrade model performance. To further explore this finding and assess the robustness of removing sentences from the front position, we conducted experiments with unguided random sentence removal (denoted as No Rule) and applied VARR+ with random position and reverse sentence order (denoted as Random and Back, respectively). As shown in Table 2, unguided random sentence removal resulted in a 25.30% decrease in performance relative to our method, highlighting the critical role of systematic configuration even after selecting sentences as the units of reduction. Additionally, Random and Back strategies result in an average 7.50% performance degradation compared to our method. This finding again indicates that earlier sentences in the reasoning path tend to be more redundant, and prioritizing their removal effectively balances rationale reduction while maintaining reasoning performance.
5.5.3 Reinitializing the Optimizer
G8K MQA CQA TQA SQA
w/o reinit 45.18 48.70 87.82 83.72 76.74
(115.91) (92.86) (32.95) (38.27) (33.89)
w/ reinit 54.98 56.95 89.56 83.45 78.19
(100.38) (91.04) (33.55) (39.17) (35.12)
Table 3: Performance Comparison before and after the reinitializing optimizer in every training epoch. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
We reinitialized the optimizer after each training epoch to stabilize training, inspired by Deng et al. (2024b). We employed the AdamW Loshchilov and Hutter (2019) optimizer, where the first and second moments are gradually updated based on current gradients, which could lead to unstable training in our framework. Specifically, after applying VARR throughout the dataset, a reduction in the rationale for some data between epochs could lead to instability in the training process. With a 19.32% performance gain achieved by reinitializing the optimizer, we reconfirm that Deng et al. (2024b)’s findings are robust across diverse datasets with complex semantic and syntactic structures in reasoning paths.
5.5.4 Varying the Warm-up Steps
Figure 4: Accuracy and Generated Token difference across various warm-up stages.
We evaluated our method against various warm-up stages on the TriviaQA and CommonsenseQA datasets. As the duration of the warm-up stages increases, the model becomes more fitted to the non-reduction dataset, which inhibits VARR’s ability to eliminate redundant sentences from the reasoning path. Consequently, as illustrated in Figure 4, longer warm-up periods result in an increase in generated tokens and a decrease in accuracy. These results suggest that 0.1 training steps provide sufficient time to inject reasoning abilities while enabling the systematic reduction of redundant reasoning sentences during the learning process. Therefore, we set 0.1 as the default setting for the warm-up steps.
5.5.5 Removal Ratio Analysis
Figure 5: Removal ratio of redundant sentences throughout the training stage. The Removal Ratio, indicated on the y-axis, is calculated by dividing the number of sentences actually removed by the maximum potential removal ratio (i.e., ${size(\{r_{i}\})/r(t)})$. The x-axis represents the index of the epochs.
In Figure 5, we analyze the actual amount of rationale sentences removed during training. We examine it by calculating the removal ratio, the proportion of actual removed sentences to the maximum potential removal sentences $r(t)$. Our analysis indicates that not all sentences designated for maximum removal range are always eliminated during the training process. Notably, a significant proportion of redundant sentences are removed in the early stages of training, with fewer sentences being removed as the model progresses through the middle to later stages, thereby stabilizing its training. A similar trend is observed across other datasets, as detailed in Appendix E.
6 Conclusion
In this work, we propose sentence-level rationale reduction framework VARR and provide empirical evidence that models trained with non-redundant rationales achieve greater efficiency. We address the lack of principled criteria for identifying redundant sentences during the training phase by developing a reduction method that not only preserves the model’s reasoning capabilities but also reduces the likelihood of generating incorrect answers. Our experiments show that VARR can efficiently handle a diverse range of tasks with fewer generated tokens, without sacrificing its accuracy. Further empirical findings confirm that sentences are appropriate reasoning units for rationale reduction, highlighting the necessity of principled criteria for robust application. This work contributes novel insights to rationale reduction research, contributing to the efficient elicitation of reasoning in language models.
Limitations
While our work provides novel insights into rationale reduction research, our experiments were primarily conducted using a relatively small large language model, constrained by computational costs. Additionally, for the same reasons, it was not feasible to test the model with datasets featuring long sequences in both queries and rationales (Reddy et al., 2024; Yu et al., 2024a). However, given that the VARR/VARR+ frameworks are designed with a systematically structured approach, we believe that they could also be effective in scaled-up settings.
Appendix A Additional Datasets Analysis
Figure 6: NLL differences across varying sizes of $\{r_{i}\}$. The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that ${r_{i}}$ are sampled from the front, random, and back indices of the full index set, respectively.
As shown in Section 3, the Negative Log-Likelihood (NLL) tends to rise with increasing size of ${r_{i}}$ in random and back configurations, compared to the front configuration, as illustrated in Figure 7 this trend is consistent across both the GSM8K and StrategyQA datasets.
Appendix B VARR Implementation
The detailed implementation of VARR+ is outlined in Algorithm 1. After the warm-up stages, each datum in the current batch is evaluated using the verbosity equation—specifically, using only Equation 5 in VARR and both Equations 5 and 9 in VARR+. As mentioned in Section 3 and 4, the process assesses the redundancy of each sentence in data starting from the first index during every single epoch. Therefore, all data can potentially shorten the length of the rationale progressively.
Algorithm 1 Training Procedure of VARR+
1:$\mathcal{D}$: Training dataset
$\mathcal{B}$: Training batch
3:$E$: Total training epochs
4:$S$: Total training steps per epoch
5:$T=E\times S$: Total training steps
6:$T_{warmup}=0.1\times T$: Warm-up steps
7:$R_{B}$: A buffer to store removed sentences
8:$r(t)$: Maximum number of removable sentences at $t$
9:$N_{t}$: Number of rationale sentences at $t$
10:$\theta$: Trained model parameters
11:for epoch = 1 to E do
12:     for step = 1 to S do
13:         Sample training batch $\mathcal{B}$ from $\mathcal{D}$
14:         $t=(epoch-1)\times S+step$
15:         if $t\leq T_{warmup}$ then
16:              pass
17:         end if
18:         for each $d\in\mathcal{B}$ do
19:              $R_{B}\leftarrow\{\}$
20:              for i = 1 to $N_{t}$ do
21:                  if $r_{i}$ satisfies Equations 5 and 9 then
22:                       remove $r_{i}$ from $d$
23:                       Add $r_{i}$ to $R_{B}$
24:                       if $|R_{B}|\geq r(t)$ then
25:                           break
26:                       end if
27:                  end if
28:              end for
29:         end for
30:         Forward pass
31:         Backward pass and update $\theta$
32:     end for
33:     Reinitialize optimizer
34:end for
Appendix C Experimental Details
For all experiments, we employ the AdamW optimizer Loshchilov and Hutter (2019), configured with a weight decay of 0.005. For the Mistral 7B model, we utilize an effective batch size of 12 with gradient accumulation set to 3, while the smaller Llama3.2 models use an effective batch size of 15. A constant learning rate of $5\times 10^{-6}$ is applied across all datasets, with bfloat16 precision. For Multiple-Choice and True/False tasks, complete sets of non-correct labels are employed to configure in-batch negatives to enhance the stability of the verbosity evaluation. To ensure a fair comparison, all baselines and methods are executed on a single A100 GPU with 80GB of memory for up to 5 epochs or 36 hours, whichever is reached first.
Appendix D Details and Statistics of Datasets
Dataset Reasoning Task Source Answer Format # TRAIN. # VALID. # TEST.
GSM8K (Cobbe et al., 2021a) Arithmetic Cobbe et al. (2021b) Number 6999 472 1319
MathQA (Amini et al., 2019a) Arithmetic Amini et al. (2019b) Multi-choice 29837 4475 2985
TriviaQA (Joshi et al., 2017) Commonsense Kim et al. (2023) Natural Language 8844 552 1659
CommonsenseQA (Talmor et al., 2019) Multi-choice Kim et al. (2023) Multi-choice 609 38 115
StrategyQA (Geva et al., 2021) Commonsense Sileo (2024) T/F 832 114 344
Table 4: Comprehensive statistics of the datasets used in our experiments are provided. GSM8K and MathQA are sourced directly from their original datasets, while the remaining datasets were obtained from Kim et al. (2023) and Sileo (2024) to access complete reasoning paths.
For our experimental analysis, we carefully selected a diverse set of five datasets used in prior works (Deng et al., 2023, 2024a; Liu et al., 2024; Yu et al., 2024b; Yin et al., 2024). To ensure explicit sentence boundaries, all datasets were processed using gpt4o-mini (OpenAI, 2024) to establish clear sentence demarcation (e.g., ‘He bikes 20*2=<<20*2=40>>40 miles each day for work So he bikes 40*5=<<40*5=200>>200 miles for work’ becomes ‘He bikes 202=<<202=40>>40 miles each day for work. So he bikes 405=<<405=200>>200 miles for work’). Table 4 comprehensively outlines each dataset, including its source and the size of the training, validation, and test samples.
Appendix E Removal Ratio Analysis on additional datasets.
Figure 7: Additional analysis of the Removal Ratio on other datasets. The Removal Ratio, indicated on the y-axis, is calculated by dividing the number of sentences actually removed by the maximum potential removal ratio (i.e., ${size(\{r_{i}\})/r(t)})$. The x-axis represents the index of the epochs.
In the GSM8K, Strategy-QA, and Math-QA datasets, similar trends are observed as in Section 5.5.5. Notably, in Strategy-QA, the Removal Ratio converges more rapidly than in other datasets. This is attributed to the dataset’s characteristics, where the sentences within the rationale predominantly list simple information rather than forming complex, interrelated structures. Consequently, redundant sentences are rapidly removed, leading to a rapid decrease in further removal activities.
Appendix F Qualitative Analysis
Table 5 illustrates the generation outputs for each method on the GSM8K test set. Explicit-CoT produces a complete reasoning path for answer generation; however, its redundancy in generating ‘needs a 3 piece place setting’ leads to hallucinations within the sentence, resulting in the incorrect final answer generation. In contrast, ICoT-SI, which removes reasoning paths at the token-level, results in a generation that lacks sufficient reasoning capabilities. Conversely, VARR+ efficiently generates key information by producing ‘\$15.00’ in a single sentence, leading to more effective and efficient generation.
Question
Avery needs to buy a 3 piece place setting (dinner & salad plate and a bowl) for her holiday dinner. She’s having 12 people over for dinner. If the dinner plates cost \$6.00 each and bowls each cost \$5.00 and the salad plates cost \$4.00, how much will she spend on place settings?
Explicit-CoT
She’s having 12 people over and needs a 3 piece place setting for each so she needs 12*3 = 36 place settings.
The dinner plates cost \$6.00 each and she needs 36 of them so that comes to 6*36 = \$216.00.
The bowl costs \$5.00 each and she needs 36 of them so that comes to 5*36 = \$180.00.
The salad plates cost \$4.00 each and she needs 36 of them so that comes to 4*36 = \$144.00.
All total, the place settings will cost \$216 for dinner plates, \$180 for bowls and \$144 for salad plates for a total of 216+180+144 = \$540.00
### 540
ICoT-SI
each so 12*6 = \$72.00 She’s buying 3 pieces of place settings (dinner plates, salad plates and bowls) for \$72.00 each so 3*72 = \$216.00
### 216
VARR+
She needs 12 place settings and each setting costs \$15.00 so she will spend 12*15 = \$180.00
### 180
Ground Truth Answer
### 180
Table 5: Qualitative Example: Generation outputs on GSM8K test set.

### v3

@@
-Verbosity-Aware Rationale Reduction: Effective Reduction of Redundant Rationale via Principled Criteria
+Verbosity-Aware Rationale Reduction: Sentence-Level Rationale Reduction for Efficient and Effective Reasoning
 Abstract
 Large Language Models (LLMs) rely on generating extensive intermediate reasoning units (e.g., tokens, sentences) to enhance final answer quality across a wide range of complex tasks.
-While generating multiple reasoning paths or iteratively refining rationales proves effective for improving performance, these approaches inevitably result in significantly higher inference costs.
-In this work, we propose a novel sentence-level rationale reduction training framework that leverages likelihood-based criteria, verbosity, to identify and remove redundant reasoning sentences.
-Unlike previous approaches that utilize token-level reduction, our sentence-level reduction framework maintains model performance while reducing generation length.
-This preserves the original reasoning abilities of LLMs and achieves an average 17.15% reduction in generation costs across various models and tasks.
+While this approach has proven effective, it inevitably increases substantial inference costs.
+Previous methods adopting token-level reduction without clear criteria result in poor performance compared to models trained with complete rationale.
+To address this challenge, we propose a novel sentence-level rationale reduction framework leveraging likelihood-based criteria, verbosity, to identify and remove redundant reasoning sentences.
+Unlike previous approaches, our method leverages verbosity to selectively remove redundant reasoning sentences while preserving reasoning capabilities.
+Our experimental results across various reasoning tasks demonstrate that our method improves performance by an average of 7.71% while reducing token generation by 19.87% compared to model trained with complete reasoning paths.
 Verbosity-Aware Rationale Reduction:
-Effective Reduction of Redundant Rationale via Principled Criteria
-Joonwon Jang1  Jaehee Kim 2  Wonbin Kweon3  Hwanjo Yu1,4,∗ POSTECH1,3,4, Seoul National University2 {kaoara, skwb4453, hwanjoyu}@postech.ac.kr 1,3,4 jaehee_kim@snu.ac.kr 2
+Sentence-Level Rationale Reduction for Efficient and Effective Reasoning
+Joonwon Jang1,5  Jaehee Kim2  Wonbin Kweon3  Seonghyeon Lee4  Hwanjo Yu1,∗ POSTECH1, Seoul National University2, University of Illinois Urbana-Champaign3 Kyungpook National University4, LG AI Research5 {kaoara, hwanjoyu}@postech.ac.kr jaehee_kim@snu.ac.kr wonbin@illinois.edu sh0416@knu.ac.kr
 1 Introduction
 Recent advances in Large Language Models (LLMs) have demonstrated remarkable reasoning capabilities comparable to human cognitive abilities (Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024).
-These models can solve complex reasoning tasks by explicitly generating extended reasoning paths.
-The generation of such paths involves producing explicit reasoning units (e.g., tokens, steps) (Yu et al., 2024b), which additionally enhances model performance through iterative prompting (Wang et al., 2023; Yao et al., 2023).
-This iterative generation of explicit reasoning paths allows the model to refine and expand its thought processes, incorporating strategic planning and ongoing cognitive generation (Xi et al., 2023; Yang et al., 2024).
-Although the extensive generation of explicit reasoning units leads to better peformance, it inherently leads to higher inference costs and increased latency (Yu et al., 2024b; Wang et al., 2024).
-Moreover, fine-tuning LLMs using a complete reasoning path does not consistently guarantee stronger performance (Yu et al., 2024b; Deng et al., 2024b; Liu et al., 2024), suggesting the need for methods that can maintain reasoning performance while reducing the generation of reasoning units.
+These works demonstrate the capability to solve complex reasoning tasks through explicitly generating extended reasoning paths.
+The generation of such paths involves producing explicit reasoning units (e.g., tokens, steps) (Yu et al., 2024b), which further enhances model performance through iterative prompting (Wang et al., 2023; Yao et al., 2023).
+Through this iterative generation of explicit reasoning paths, the model refines and expands its thought processes while incorporating strategic planning and continuous cognitive generation (Xi et al., 2023; Yang et al., 2024).
+While the extensive generation of explicit reasoning units leads to improved performance, it inevitably results in higher inference costs and increased latency (Yu et al., 2024b; Wang et al., 2024).
+Furthermore, fine-tuning LLMs with complete reasoning paths does not necessarily guarantee enhanced performance (Yu et al., 2024b; Deng et al., 2024b; Liu et al., 2024), indicating the necessity for methods that maintain reasoning capabilities while reducing the generation of reasoning units.
 Despite this apparent requirement, it remains underexplored how to maintain LLM reasoning capabilities while reducing intermediate reasoning paths across diverse tasks.
-Previous methods (Yu et al., 2024b) have drawn inspiration from human cognitive processes to address the aforementioned limitations.
-Some studies have proposed training pipelines that utilize augmented datasets, iteratively generated by foundation LLMs, to fine-tune subsequent LLMs (Yu et al., 2024b; Liu et al., 2024).
-However, they are inherently vulnerable as they significantly depend on the generative capabilities of LLMs.
-In response, other works have directly trained LLMs without augmenting datasets that facilitate implicit reasoning before answer generation.
+Previous methods primarily focused on reducing reasoning paths from two distinct perspectives.
+Some studies have proposed training pipelines that leverage augmented datasets, iteratively generated by foundation LLMs, to fine-tune subsequent LLMs (Yu et al., 2024b; Liu et al., 2024).
+However, these approaches remain inherently vulnerable due to their significant dependence on the generative capabilities of LLMs.
+In response, other works have focused on directly training LLMs without dataset augmentation to reduce explicit reasoning paths.
 Deng et al.
@@
 Deng et al.
-(2024b) selected tokens as the reasoning unit for reduction and proposed a heuristic method to internalize explicit intermediate rationale tokens.
-They progressively eliminate tokens from the beginning of the full rationale during the Chain of Thought (CoT) fine-tuning.
-However, these methods present a fundamental limitation as they lack sufficient justification for selecting tokens over more linguistically natural units (e.g., sentences) for reduction (Table 1), and they fail to provide principled criteria for the removal process.
-Furthermore, their evaluation has been predominantly conducted on synthetic arithmetic reasoning tasks, which presents limitations for real-world applications.
-To address these limitations, we propose a novel training method for maintaining LLM reasoning performance while systematically eliminating redundant reasoning units within the reasoning process.
-Our method leverages sentences as fundamental reduction units, establishing more linguistically coherent boundaries compared to token-level approaches.
-We first empirically demonstrate that rationale sentences in early reasoning stages can introduce redundancy in LLM’s answer generation process.
-Drawing inspiration from Dong et al.
-(2023), we introduce the concept of ‘verbosity’ and develop a CoT fine-tuning method that removes unnecessary rationale sentences, thereby reducing rationale generation while preserving reasoning capabilities.
-The primary contributions of this work are as follows:
-  1. 1.
-    We propose a novel method for sentence-level reasoning path reduction in LLMs.
-  2. 2.
-    We demonstrate that principled criteria can effectively identify redundancies and reduce intermediate token generation efficiently.
-  3. 3.
-    We validate our method’s effectiveness and robustness through extensive experiments across various LLMs and ablation studies.
+(2024b) adopted tokens as the reasoning unit for reduction and proposed a heuristic method to internalize explicit intermediate rationale tokens while Hao et al.
+(2024) compressed complete rationales into the predefined number of hidden states of tokens.
+However, their methods present a fundamental limitation as they lack sufficient justification for selecting tokens over more linguistically natural units (e.g., sentences) for reduction (Table 1), and they fail to provide principled criteria for the removal process.
+Moreover, their evaluation has primarily focused on synthetic arithmetic reasoning tasks, limiting their applicability to real-world scenarios.
+To address these limitations, we propose a novel training method that maintains LLM reasoning performance while systematically reducing redundant reasoning units within the reasoning process.
+Our method adopts sentences as fundamental reduction units, establishing more linguistically meaningful boundaries compared to token-level approaches.
+Through empirical analysis, we demonstrate that sentences in early rationale steps can introduce redundancy in the LLM’s answer generation process.
+Inspired by Dong et al.
+(2023), we introduce the concept of ‘verbosity’, a likelihood-based criteria, to identify redundant reasoning sentences.
+By incorporating verbosity identification into the training process, the model excludes redundant reasoning sentences, thereby reducing intermediate token generation.
+Finally, we demonstrate our method’s effectiveness and generalizability across various real-world reasoning datasets, showing our method improves performance by an average of 7.71% while reducing token generation by 19.87% across various LLMs, and through systematic ablation studies, we analyze the contribution of each proposed component.
 2 Related Works
 2.1 Performance-Cost Tradeoffs in Reasoning Path Generation
-Recent research has demonstrated the importance of generating iterative and refined reasoning paths to improve the reasoning capabilities of models, albeit at increased computational costs (Wang et al., 2023; Yao et al., 2023; Radha et al., 2024; Wang et al., 2024; Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024).
-Self-Consistency (Wang et al., 2023) improves reasoning accuracy by generating more than 20 distinct reasoning paths and selecting the most consistent answer using ensemble methods, resulting in substantial computational overhead.
-Similarly, Tree of Thoughts (ToT) (Yao et al., 2023) expands the exploration of reasoning paths using a tree-structured method, while Strategic Chain of Thought (SCoT) (Wang et al., 2024) utilizes an Inner Dialogue Agent to adaptively explore the diverse reasoning path.
-Concurrently, Self-Refine (Madaan et al., 2024) utilizes an iterative framework that incorporates feedback to refine responses and enhance reasoning accuracy, necessitating multiple forward passes through the model.
-Additionally, the Reflexion-based framework (Shinn et al., 2024; Kumar et al., 2024) guides models to iteratively generate relevant reasoning paths through reflection for accurate answer generation.
-Given the necessity of iterative generation and refinement of reasoning paths to achieve optimal performance, they inherently increase inference costs and latency.
-Therefore, it is essential to investigate methods for efficiently generating these paths.
+Recent research has demonstrated the critical role of generating iterative and refined reasoning paths in enhancing model reasoning capabilities, albeit at increased computational costs (Wang et al., 2023; Yao et al., 2023; Radha et al., 2024; Wang et al., 2024; Madaan et al., 2024; Shinn et al., 2024; Kumar et al., 2024).
+Self-Consistency (Wang et al., 2023), Tree of Thoughts (ToT) (Yao et al., 2023), and Strategic Chain of Thought (SCoT) (Wang et al., 2024) improve reasoning accuracy through ensemble-based path selection, tree-structured exploration, and adaptive reasoning with an Inner Dialogue Agent, where each approach requires iterative reasoning path generation, resulting in substantial computational overhead.
+Concurrently Self-Refine (Madaan et al., 2024) and Reflexion-based framework (Shinn et al., 2024; Kumar et al., 2024) enhance reasoning abilities through iterative feedback-based refinement and reflective path generation, respectively, though both require multiple forward passes through the model.
+While the iterative generation and refinement of reasoning paths are essential for achieving optimal performance, they inherently increase inference costs and latency.
+Therefore, it is crucial to investigate methods for efficiently generating these paths.
 2.2 Reasoning Path Reduction
-To address the computational costs associated with extensive reasoning paths generation, some lines of work (Yu et al., 2024b; Liu et al., 2024) have focused on generating augmented datasets with varying rationale lengths to shorten the generation of reasoning paths.
+To address the computational costs associated with extensive reasoning paths generation, some lines of work (Yu et al., 2024b; Liu et al., 2024) have focused on generating augmented datasets with varying rationale lengths to reduce the generation of reasoning paths.
 Yu et al.
-(2024b) guide model to generate multiple reasoning paths using Self-Consistency to construct datasets, subsequently fine-tuning the model to generate direct answers without intermediate reasoning steps.
+(2024b) employs Self-Consistency to generate multiple reasoning paths for dataset augmentation, then fine-tunes the model to produce direct answers.
 Liu et al.
-(2024) developed a heuristic approach to merge reasoning steps and iteratively trained the model to produce shorter reasoning paths, which are subsequently incorporated into the progressive training phase.
-While these approaches demonstrate empirical effectiveness, they present two fundamental limitations: (1) their heavy dependence on LLM generation capabilities introduces inherent instability, and (2) their objective of reducing reasoning path necessitates the paradoxical creation of datasets requiring extensive reasoning path generation.
-To address these weaknesses, another line of work (Deng et al., 2023, 2024b) has focused on directly training LLM without augmented datasets.
-Implicit-CoT (Deng et al., 2023) utilizes a multi-model framework where an emulator model is trained to predict the teacher’s token-level hidden states, and a student model uses these predicted states to generate answers.
-ICoT-SI (Deng et al., 2024b) identifies tokens as reduction units, proposing a method to internalize explicit intermediate rationale tokens by progressively removing them from the beginning of the reasoning path within the CoT fine-tuning process.
-However, these methods have shown limited generalization across diverse datasets as they have only been validated on simple arithmetic reasoning tasks, such as multiplication problems.
+(2024) developed a heuristic approach to merge reasoning steps and iteratively trained the model to produce shorter reasoning paths, which are then integrated into the progressive training phase.
+While these approaches demonstrate empirical effectiveness, they exhibit two fundamental limitations: (1) their substantial dependence on LLM generation capabilities introduces inherent instability, and (2) their objective of reducing reasoning paths necessitates the paradoxical creation of datasets requiring extensive reasoning path generation.
+To address these weaknesses, another line of work (Deng et al., 2023, 2024b; Hao et al., 2024) has focused on directly training LLMs without augmented datasets.
+Implicit-CoT (Deng et al., 2023) implements a multi-model framework where an emulator model is trained to predict the teacher’s token-level hidden states, and a student model leverages these predicted states to generate answers.
+ICoT-SI (Deng et al., 2024b) identifies tokens as reduction units, proposing a method to internalize explicit intermediate rationale tokens by progressively eliminating them from the beginning of the reasoning path within the CoT fine-tuning process.
+However, these methods demonstrate limited generalization across diverse datasets as they have been validated exclusively on simple arithmetic reasoning tasks, such as multiplication problems.
 This limitation raises concerns about their applicability to real-world scenarios where rationales are expressed in natural language.
-Moreover, they do not justify their criteria for token-level removal nor consider whether more linguistically natural units (e.g., sentences) might serve as more effective reduction units.
-Specifically, the token-level removal approach may inadvertently eliminate critical information necessary for answer generation or distort the semantic information of the sentence, while requiring repeated removal of redundant reasoning steps.
-Motivated by the computational costs of the extensive reasoning path generation and the limitations of existing reduction approaches, we investigate the redundancy of various sentence positions for potential removal and propose a novel method with principled criteria that can efficiently reduce reasoning paths while maintaining their effectiveness.
-3 Early Stage Rationales are Redundant
+Furthermore, they do not explore the adoption of principled criteria and linguistically natural units (e.g., sentences).
+Specifically, the token-level reduction approach may eliminate critical information necessary for answer generation or distort the semantic information of the sentence.
+Motivated by these limitations in existing reduction approaches, we examine the redundancy of various sentence positions for potential elimination and propose a novel method with principled criteria that can effectively reduce reasoning paths while maintaining their efficacy.
+3 Early Step Rationales are Redundant
 Figure 1: NLL differences across varying sizes of $\{r_{i}\}$.
-The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that $\{r_{i}\}$ is sampled from the front, random, and back indices of the full index set, respectively.
+The ‘original’ represents the NLL with the complete rationale, while ‘front’, ‘random’, and ‘back’ indicate that $\{r_{i}\}$ is sampled from the front, random, and back indices of the full index set, respectively.
 3.1 Quanitifying the Redundancy
 Before delving into the method, we first investigate which positions within the rationale sentences should be selected for reduction.
-When the likelihood of the answer remains unchanged after a sentence is removed from the full rationales, it indicates that the sentence may be redundant for the reasoning process.
-To quantify the redundancy of a sentence, we compute the negative log-likelihood (NLL) for answer $y$ after the sentence reduction as follows:
+When the likelihood of the answer remains unchanged after removing a sentence from the full rationales, this indicates that the sentence may be redundant in the reasoning process.
+To quantify the redundancy of a sentence, we compute the negative log-likelihood (NLL) for answer $y$ after sentence reduction as follows:
 $\begin{split}\text{NLL}&=-\log p_{\theta}(y|R^{\prime},x),\\ \text{where}\ R^{\prime}&=R\setminus\{r_{i}\}_{i\in S},\ S\subseteq I.\end{split}$ (1)
@@
 For simplicity, we use $\{r_{i}\}$ without the subset index notation throughout the rest of the paper.
-3.2 Redundancy of leading sentences
-We conduct a pilot study to empirically demonstrate the redundancy of the leading sentences within the rationales.
-By varying the size of $\{r_{i}\}$ from 1 to 4, we calculate the NLL of Mistral 7B (Jiang et al., 2023) across different reasoning datasets.
-To explore more general patterns of sentence removal, we employ a stochastic strategy in our experimental setup.
-We compare three configurations for sampling $\{r_{i}\}$: front, random, and back.
-In the front configuration, initial sentences are assigned a higher probability when forming $R^{\prime}$, whereas random and back configurations use uniform and progressively increasing probabilities, respectively [^1] For generating $R^{\prime}$, we assign probabilities $p_{k}=\frac{N-k+1}{\sum_{i=1}^{N}i}$ (front), $\frac{1}{N}$ (random), and $\frac{k}{\sum_{i=1}^{N}i}$ (back) where k=1,…,N denotes sentence position..
-Additionally, we computed the NLL for full rationale sentences (i.e., $-\log p_{\theta}(y|R,x)$) as a baseline to evaluate the impact of reduction.
-As depicted in Figure 1, the NLL difference between the full and reduced rationales shows a marginal likelihood difference in the front ($<$) configuration.
-In contrast, removing sentences randomly ($\diamond$) or from the back ($>$) results in higher NLL as the removed sentences increase, highlighting the importance of the selection of removal position in the reasoning and answer prediction process (for additional data, see Appendix A).
+3.2 Redundancy of Early Reasoning Sentences
+We performed a pilot study to empirically demonstrate the redundancy of leading sentences within the rationales by analyzing the NLL of Mistral 7B (Jiang et al., 2023) across diverse reasoning datasets.
+Specifically, we varied the size of $\{r_{i}\}$ from 1 to 4 and investigated general patterns of sentence removal using a stochastic approach.
+To compare different sentence selection configurations, we considered three sampling methods for $\{r_{i}\}$: front, where initial sentences were prioritized; random, with uniform probabilities; and back, where probabilities progressively increased for later sentences [^1] For generating $R^{\prime}$, we assign probabilities $p_{k}=\frac{N-k+1}{\sum_{i=1}^{N}i}$ (front), $\frac{1}{N}$ (random), and $\frac{k}{\sum_{i=1}^{N}i}$ (back) where k=1,…,N denotes sentence position..
+Additionally, we computed the NLL for complete rationale sentences (i.e., $-\log p_{\theta}(y|R,x)$) as a baseline to evaluate the impact of reduction.
+As illustrated in Figure 1, the front ($\triangleleft$) configuration shows only marginal NLL differences relative to complete rationale sentences.
+In contrast, removing sentences randomly ($\diamond$) or from the back ($\triangleright$) results in higher NLL as the removed sentences increase, highlighting the importance of the selection of a candidate rationale position strategy for removal in the reasoning and answer prediction process (for additional analysis, see Appendix A).
 4 Verbosity-Aware Rationale Reduction
@@
 Initially, we select a candidate sentence from the beginning of the rationale (Section 3).
-After removing the candidate sentence, we evaluate $\textit{verbosity}(y_{g})$ and $\textit{verbosity}(y_{w})$ using Equations (5) and (9).
+After selecting the candidate sentence, we evaluate Equations (5) and (9) by calculating $\textit{verbosity}(y_{g})$ and $\textit{verbosity}(y_{w})$.
 If the candidate sentence meets the verbosity evaluation criteria, it is excluded in subsequent training steps.
 The model then proceeds with training, where the redundant sentence is excluded from the rationale.
-Based on the above observation, we propose the Verbosity-Aware Rationale Reduction (VARR) framework.
+Based on these observations, we propose the Verbosity-Aware Rationale Reduction (VARR) framework.
 In Section 4.1, we introduce the concept of ‘verbosity’ as a principled criterion for identifying redundant reasoning sentences.
-Section 4.2 describes the integration of verbosity in the reduction process during CoT training.
-Subsequently, Section 4.3 presents additional verbosity term by incorporating the wrong answer to improve robustness.
-Finally, Section 4.4 illustrates the overall VARR framework.
-4.1 Verbosity
-To quantify the redundancy of the sentences for removal, we first introduce a foundational concept ‘verbosity’.
-Given an input $x$, full rationale $R$, and a reduced rationale $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, we measure the verbosity of a sentence $r_{i}$ on $y$ by measuring the difference in Kullback–Leibler divergence (KL-divergence) as follows:
+Section 4.2 elaborates on how we integrate verbosity into the reduction process during CoT training.
+In Section 4.3, we extend the verbosity term by incorporating incorrect answers to enhance robustness.
+Finally, Section 4.4 presents the comprehensive VARR framework.
+4.1 Verbosity as Principled Criterion
+To quantify the redundancy of sentences for potential removal, we introduce the fundamental concept ‘verbosity’.
+Given an input $x$, full rationale $R$, and a reduced rationale $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, we quantify the verbosity of a sentence $r_{i}$ on $y$ by computing the difference in Kullback–Leibler divergence (KL-divergence) as follows:
 $\begin{split}\textit{verbosity}(y)&=D_{KL}(q(y|x)\parallel p_{\theta}(y|R,x))\\ &\hskip 15.00002pt-D_{KL}(q(y|x)\parallel p_{\theta}(y|R^{\prime},x)),\end{split}$ (2)
 where $q(y|x)$ is the ground truth distribution.
-The verbosity$(y)$ highlights the informational contribution or redundancy of a rationale sentence $r_{i}$ on answer $y$.
-Since $q(y|x)$ is the form of the one-hot vector (i.e., Dirac delta function), the verbosity$(y)$ can be computed as the log-likelihood ratio between $R$ and $R^{\prime}$ as follows:
-$\begin{split}\textit{verbosity}(y_{g})&=\left[H_{q}(p_{\theta}(y|R,x))-H(q(y|x))\right]\\ &\hskip 15.00002pt-\left[H_{q}(p_{\theta}(y|R^{\prime},x))-H(q(y|x))\right]\\ &=E_{q}[-\log P(y|R,x)]\\ &\hskip 15.00002pt+E_{q}[\log P(y|R^{\prime},x)]\\ &=\log\left(\frac{P(y_{g}|R^{\prime},x)}{P(y_{g}|R,x)}\right),\end{split}$ (3)
-where $y_{g}$ is the ground truth answer (i.e., $q(y_{g}|x)=1$).
-$H_{q}(\cdot)$ and $H(\cdot)$ denote the cross-entropy and the entropy, respectively.
+The verbosity$(y)$ measures the informational contribution or redundancy of a rationale sentence $r_{i}$ with respect to answer $y$.
+Since $q(y|x)$ is the form of the one-hot vector (i.e., Dirac delta function), we can express the verbosity$(y)$ as the log-likelihood ratio between $R$ and $R^{\prime}$ as follows:
+$\begin{split}\textit{verbosity}(y_{g})&=\left[H_{q}(p_{\theta}(y|R,x))-H(q(y|x))\right]\\ &\hskip 6.99997pt-\left[H_{q}(p_{\theta}(y|R^{\prime},x))-H(q(y|x))\right]\\ &=E_{q}[-\log p_{\theta}(y|R,x)]\\ &\hskip 6.99997pt+E_{q}[\log p_{\theta}(y|R^{\prime},x)]\\ &=\log\left(\frac{p_{\theta}(y_{g}|R^{\prime},x)}{p_{\theta}(y_{g}|R,x)}\right),\end{split}$ (3)
+where $y_{g}$ denotes the ground truth answer (i.e., $q(y_{g}|x)=1$).
+$H_{q}(\cdot)$ and $H(\cdot)$ denote the cross-entropy and the entropy, respectively [^2] For the sake of explainability, we assume each expression’s $y$ is represented by a single token..
 Intuitvely, a higher value of $\textit{verbosity}(y_{g})$ implies that the likelihood of the model generating the ground truth answer increases after removing $r_{i}$, indicating that its removal is beneficial.
 4.2 Verbosity Identification in CoT Training
-Given an input sequence, the CoT training (Nye et al., 2021) aims to train LLMs to generate full rationale, followed by the ground truth answer with the following loss:
+Given an input sequence, CoT training (Nye et al., 2021) aims to train LLMs to generate complete rationale, followed by the ground truth answer:
 $-\log p_{\theta}(y_{g},R|x).$ (4)
-In each training step $t$, each sentence $r_{i}$ within $R$ is assessed with the following the equation:
+During each training step $t$, we evaluate each sentence $r_{i}$ within $R$ using the following criterion:
 $\textit{verbosity}(y_{g})\geq 0.$ (5)
 Here, we sequentially select $r_{i}$ starting from the first sentence and construct $R^{\prime}=\{r_{j}\}_{j\in I\setminus\{i\}}$, based on our analysis in Section 3, which indicates that early-stage rationales are more likely to be redundant.
-A non-positive verbosity score indicates that removing $r_{i}$ from $R$ impairs the model’s performance, thereby identifying the sentence as essential and retaining it in the rationale.
+When the verbosity score is non-positive, it indicates that removing $r_{i}$ from $R$ would impair the model’s performance, thus identifying the sentence as essential and preserving it in the rationale.
+Detailed training procedures will be described in Section 4.4.
 4.3 Contrasting with Wrong Answer
-Inspired by the miscalibrated log-likelihoods between accepted and rejected responses during standalone Supervised Fine-Tuning (SFT) in alignment learning Rafailov et al.
+Inspired by the miscalibrated log-likelihoods between accepted and rejected responses in standalone Supervised Fine-Tuning (SFT) for alignment learning Rafailov et al.
 (2024); Azar et al.
 (2024); Hong et al.
-(2024), we also assess whether the reduced rationales $R^{\prime}$ result in inaccuracy answer generation by introducing a wrong answer $y_{w}$.
-Instead of employing the correct answer distribution $q(y|x)$ in Equation (2), we initiate our formula with the wrong answer distribution $q^{\prime}$ (i.e., $1-q(y|x)$, normalized sum to 1) Through algebraic manipulation, we can obtain:
+(2024), we examine whether reduced rationales $R^{\prime}$ lead to inaccurate answer generation by incorporating a wrong answer $y_{w}$.
+Instead of employing the correct answer distribution $q(y|x)$ in Equation (2), we initiate our formula with the wrong answer distribution $q^{\prime}$ (i.e., $1-q(y|x)$, normalized sum to 1).
+Through algebraic manipulation, we derive:
 $\begin{split}&D_{KL}(q^{\prime}(y|x)\parallel p_{\theta}(y|R,x))\\ &\hskip 10.00002pt-D_{KL}(q^{\prime}(y|x)\parallel p_{\theta}(y|R^{\prime},x))\\ &=\left[H_{w}(p_{\theta}(y|R,x))-H(q^{\prime}(y|x))\right]\\ &\hskip 10.00002pt-\left[H_{w}(p_{\theta}(y|R^{\prime},x))-H(q^{\prime}(y|x))\right],\end{split}$ (6)
@@
 $\begin{split}&E_{w}[-\log p_{\theta}(y|R,x)]+E_{w}[\log p_{\theta}(y|R^{\prime},x)]\\ &={E}_{w}\Big{[}\log\frac{p_{\theta}(y|R^{\prime},x)}{p_{\theta}(y|R,x)}\Big{]}\\ &\approx\frac{1}{K}\sum_{k=1}^{K}\log\frac{p_{\theta}(y_{w}^{(k)}|R^{\prime},x)}{p_{\theta}(y_{w}^{(k)}|R,x)},\end{split}$ (7)
-where $\{y_{w}^{(k)}\}_{k\in\left[K\right]}$ is sampled from the in-batch negatives depending on the dataset. Consequently, the $\textit{verbosity}(y_{w})$ is computed as:
+where $\{y_{w}^{(k)}\}_{k\in\left[K\right]}$ is sampled from the in-batch negatives depending on the dataset. Consequently, $\textit{verbosity}(y_{w})$ is computed as:
 $\textit{verbosity}(y_{w})=\frac{1}{K}\sum_{k=1}^{K}\log\frac{p_{\theta}(y_{w}^{(k)}|R^{\prime},x)}{p_{\theta}(y_{w}^{(k)}|R,x)}.$ (8)
-Since computational constraints necessitate sampling incorrect answers to calculate $\textit{verbosity}(y_{w})$, we contrast $\textit{verbosity}(y_{w})$ with $\textit{verbosity}(y_{g})$ rather than solely using $\textit{verbosity}(y_{w})$.
-This ensures that the likelihood of incorrect answers remains lower than that of ground truth answers, even after the removal of $r_{i}$ as follows:
+Since computational constraints necessitate sampling incorrect answers to calculate $\textit{verbosity}(y_{w})$, we evaluate the effectiveness of removal by comparing $\textit{verbosity}(y_{w})$ against $\textit{verbosity}(y_{g})$ rather than solely using $\textit{verbosity}(y_{w})$ as follows:
 $\textit{verbosity}(y_{w})-\textit{verbosity}(y_{g})\leq 0.$ (9)
-A smaller $\textit{verbosity}(y_{w})$ value compared to $\textit{verbosity}(y_{g})$ indicates that $R^{\prime}$ increases the likelihood of generating incorrect answers over the ground truth answer after removing $r_{i}$, making it beneficial to remove.
+When both conditions $\textit{verbosity}(y_{g})\geq 0$ and $\textit{verbosity}(y_{w})-\textit{verbosity}(y_{g})\leq 0$ are satisfied, it indicates that removing $r_{i}$ from $R$ not only improves the model’s performance but also increases its preference for the ground truth answer over incorrect answers, supporting the removal of $r_{i}$.
 4.4 CoT Training with Rationale Reduction
-In our framework, the model is trained using Equation (4) for predefined warm-up steps to inject its reasoning capabilities.
-After the warm-up steps, at each training step $t$, each sentence $r_{i}$, starting from the first index sentence in $R$, is evaluated against either Equation (5) alone (denoted as VARR) or both Equations (5) and (9) (denoted as VARR+).
-Sentences satisfying these respective criteria are selected for removal and are not included in subsequent training steps.
-The maximum removable number of sentences at each training step $t$ is determined based on a linear schedule, adopting Deng et al.
-(2024b)’s setting as follows:
+In our framework, the model is trained using Equation (4) for predefined warm-up stage to inject its reasoning capabilities.
+Subsequently, at each training step $t$, we evaluate each sentence $r_{i}$ sequentially from the first sentence in $R$, using either Equation (5) alone (denoted as VARR) or the combination of Equations (5) and (9) (denoted as VARR+).
+Sentences that satisfy these respective criteria are identified for removal and excluded from subsequent training steps.
+The maximum removable number of sentences at each training step $t$ is determined based on a linear schedule, adopting ICoT-SI (Deng et al., 2024b)’s setting as follows:
 $r(t)=\lfloor N_{t}\cdot(t/T)\rfloor,$ (10)
 where $T$ represents the total number of training steps, $N_{t}$ is the total number of rationale sentences at step $t$, and $r(t)$ indicates the maximum number of sentences that can be removed at that step.
+Note that unlike ICoT-SI, which enforcely removes a predefined number of tokens during training, our method preserves the essential reasoning steps by employing principled removal criteria.
 5 Experiments
 5.1 Training Configuration
-Figure 3: Performance for 5 different reasoning tasks.
-Acc refers to Accuracy while Tokens denotes the Generated Tokens.
-Bar plots illustrate the accuracy of each method on datasets, while line plots show the average generated tokens by each method across the datasets.
-The y-axis represents the range of each metric, and the x-axis lists the datasets evaluated.
+Figure 3: Pareto plot of accuracy versus the number of generated tokens.
+The gray dotted lines connect the Pareto frontiers of the baselines, and our VARR (or VARR+) consistently outperforms the pareto frontiers across all subplots.
+While ICoT-SI and Coconut substantially trade-off accuracy for efficiency, VARR/VARR+ maintains high accuracy while reducing generated tokens, demonstrating its superior efficiency-performance balance.
 Datasets
 We conducted experiments across two categories to provide a comprehensive evaluation of VARR’s versatility and effectiveness, in contrast to prior research that predominantly focuses on simple arithmetic tasks like multi-digit multiplication (Deng et al., 2023, 2024b).
-Initially, we evaluate with arithmetic reasoning tasks, including datasets like GSM8K (G8K; Cobbe et al.
-2021a) and MathQA (MQA; Amini et al.
-2019a).
+Initially, we evaluate with arithmetic reasoning tasks, including datasets like MathQA (MQA; Amini et al.
+2019a) and GSM8K (G8K; Cobbe et al.
+2021a).
 We also examine the performance of our method on commonsense reasoning tasks, employing datasets including CommonsenseQA (CQA; Talmor et al.
-2019), TrivaiQA (TQA; Joshi et al.
+2019), TriviaQA (TQA; Joshi et al.
 2017), and StrategyQA (SQA; Geva et al.
 2021).
+Note that unlike previous works (Deng et al., 2023, 2024b), we do not synthesize training data (especially the intermediate steps) to validate the generalizability and applicability of our method.
 Models
@@
 In addition, an optimizer is reinitialized at the beginning of each epoch to stabilize the model training inspired by Deng et al.
-(2024b)’s setting, with its effects described in Section 5.5.3.
-All baselines and our proposed method are trained for 5 epochs to ensure fairness in comparison.
-Detailed implementation of VARR/VARR+ can be found in the Appendix B.
+(2024b)’s setting, with its effects described in Appendix E.
+All methods are trained for 5 epochs for fair comparison, with detailed implementation of VARR/VARR+ provided in Appendix B.
 5.2 Baselines
-We compare our method to the following baselines: Explicit-CoT (Nye et al., 2021), where the model is finetuned with explicit chain-of-thought reasoning, and ICoT-SI (Deng et al., 2024b), where the model is fine-tuned using a linear token elimination schedule.
-It is important to note that after removing intermediate reasoning paths, the model’s performance should be competitive with Explicit-CoT.
-Consequently, ICoT-SI serves as our primary baseline.
-We exclude Implicit-CoT Deng et al.
-(2023) from our baselines due to its significant memory requirements, requiring the training of three models on a single GPU.
-Furthermore, according to Deng et al.
-(2024b), this method underperformed compared to ICoT-SI.
+We compared our method against the following baselines: Explicit-CoT (Nye et al., 2021), where the model is finetuned with explicit chain-of-thought reasoning; ICoT-SI (Deng et al., 2024b), where the model is fine-tuned using a linear token elimination schedule; and Coconut (Hao et al., 2024), where the model is fined-tuned to compress rationales into a predefined number of token hidden states.
+We excluded Implicit-CoT Deng et al.
+(2023) from our evaluation due to its substantial computational demands, specifically requiring three models to be trained simultaneously on a single GPU.
+Moreover, Deng et al.
+(2024b) demonstrated that this method achieves a lower performance compared to ICoT-SI.
+Given that both our baselines and VARR aim to maintain Explicit-CoT’s performance while reducing the number of generated tokens, we establish Explicit-CoT’s performance as our primary baseline for comparison.
 All baselines were trained on a single A100-80GB GPU, and detailed training configurations for each method are provided in Appendix C.
@@
 In Figure 3, we present the results for each reasoning task across different models.
-Our method demonstrates consistent performance improvements or maintains comparable results across most benchmarks, while also achieving more efficient reasoning path generation.
-Notably, we observed significant improvements in average accuracy with increments of 2.59%, 6.27%, and 14.27% for Mistral 7B, Llama3.2 1B, and Llama3.2 3B (VARR+), respectively, compared to Explicit CoT, while efficiency improved by 19.35%, 17.24%, and 14.88% for these models.
-These findings indicate that some reasoning sentences are redundant and can even disturb answer generation, highlighting that effective reasoning capabilities can be learned through selective utilization of essential reasoning units.
-For ICoT-SI, the model exhibited performance degradation after token removal training.
-Specifically, ICoT-SI shows an average performance degradation of 18.93% compared to Explicit CoT.
-Furthermore, while the Mistral 7B model achieved a 50.26% accuracy on the GSM8K using a large augmented dataset with succinctly generated reasoning paths (trained with the 385K-size dataset used in the original paper (Deng et al., 2024b)), its performance dropped significantly to 28.27% when trained on the standard GSM8K dataset (trained with the 7K-size dataset).
-This indicates that while ICoT-SI may excessively reduce the length of rationales in reasoning tasks, identifying the appropriate reasoning unit and applying principled criteria are crucial for maintaining performance in real-world application problems.
-Qualitative results across baselines are presented in Appendix F, and a detailed analysis of the reasoning unit will be discussed in Section 5.5.1.
-Moreover, introducing a wrong answer in VARR led to performance improvements in most datasets.
-While solely using the verbosity($y_{g}$) enhanced efficient generation, it is notable that verbosity($y_{w}$) constraints reduced rationales to increase the likelihood of generating incorrect answers, which contributes to the robustness and stability of the training process.
+VARR/VARR+ achieves comparable or superior performance compared to Explicit-Cot across most datasets while reducing the average token generation.
+Specifically, VARR+ significantly increases performance by an average of 7.71% across all datasets and models, while improving efficiency by reducing token generation by 19.37% on average.
+It is noteworthy that these findings contrast with ICoT-SI, demonstrating that performance can be improved while reducing the number of generated tokens.
+This suggests that existing reasoning data contains unnecessary reasoning sentences that may harm performance.
+Furthermore, effective reasoning can be achieved by selectively removing sentences based on appropriate criteria.
+However, the baselines exhibit performance degradation compared to Explicit-CoT.
+For ICoT-SI, we observe an average performance decline of 21.98%, while Coconut shows a degradation of 25.20%, demonstrating their imbalanced trade-off in efficiency.
+These results suggest that heuristic reasoning reduction approaches do not effectively induce implicit reasoning within the model as discussed by ICoT-SI and Coconut.
+This indicates that identifying and retaining appropriate reasoning units through principled criteria is crucial for maintaining performance in practical applications.
+A more detailed discussion of the impact of the choice of reduction unit and criterion is provided in Section 5.5.1.
+Furthermore, qualitative analysis (refer to Appendix J) confirms that ICoT-SI and Coconut fail to generate valid reasoning paths for answer generation, while VARR+ produces concise, yet effective reasoning paths that lead to correct answers.
+Additionally, the incorporation of incorrect answers in VARR+ resulted in performance improvements across most datasets.
+While VARR alone enhanced generation efficiency, VARR+ effectively preserves rationales that help calibrate the probability distribution between correct and incorrect answers, contributing to improved training robustness and stability.
 5.5 Ablation Studies
@@
 5.5.1 Identifying Appropriate Units for Removal
-G8K MQA CQA TQA SQA
-Exp-CoT 55.26 55.84 84.33 82.94 74.70
+MQA G8K CQA TQA SQA
+Exp-CoT 55.84 55.26 84.33 82.94 74.70
 (105.02) (122.54) (42.84) (63.46) (46.47)
-ICoT-SI 28.27 35.84 67.82 77.09 61.33
+ICoT-SI 35.84 28.27 67.82 77.09 61.33
 (113.55) (104.41) (6.0) (22.03) (22.86)
-VARR-Tok 47.53 46.79 82.60 67.14 71.22
+VARR-Tok 46.79 47.53 82.60 67.14 71.22
 (90.39) (94.16) (25.07) (38.10) (31.92)
-VARR-Sent 54.98 56.95 89.56 83.45 78.19
+VARR-Sent 56.95 54.98 89.56 83.45 78.19
 (91.04) (100.38) (33.55) (39.17) (35.12)
-Table 1: Analysis across various reasoning reduction units and the application of principled criteria. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
-In this section, we experimentally demonstrate why the sentences are suitable units for reduction reasoning units.
-Unlike ICoT-SI which lacks specific criteria for token removal, we apply VARR+ at the token level (denoted as VARR-Tok) to assess whether tokens can effectively function as reduction units in reasoning processes.
-As depicted in Table 1, applying principled criteria at the token level results in an average performance gain of 26.95%, indicating that the principled criteria contribute to robust performance.
-Moreover, extending the reasoning units from tokens to sentences (denoted as VARR-Sent) leads to a further performance increase of 21.83% over VARR-Tok, highlighting the importance of selecting the appropriate reasoning unit for effective reasoning path reduction.
-These findings demonstrate the importance of selecting appropriate reasoning units, with sentences serving as linguistically natural and effective boundaries for the removal process.
+Table 1: Analysis across various reasoning reduction units and the application of principled criteria.
+Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
+In this section, we empirically investigate the necessity of our criteria and demonstrate why sentences are more effective than tokens as reasoning reduction units.
+While ICoT-SI removes tokens without specific criteria, we first apply VARR+ at the token level (denoted as VARR-Tok) to assess whether tokens can serve as effective reduction units when combined with our criteria.
+As shown in Table 1, VARR+ applied at the token level achieves an average performance gain of 24.74% compared to ICoT-SI, demonstrating that our principled criteria contribute to robust performance.
+Furthermore, expanding the reduction units from tokens to sentences (denoted as VARR-Sent) yields an additional performance gain of 15.98% over VARR-Tok.
+These findings highlight that sentences provide natural and effective boundaries for the reduction process.
 5.5.2 Analyzing the Impact of Sentence Position on Removal Efficacy
-G8K MQA CQA TQA SQA
-Exp-CoT 55.26 55.84 84.33 82.94 74.70
+MQA G8K CQA TQA SQA
+Exp-CoT 55.84 55.26 84.33 82.94 74.70
 (105.02) (122.54) (42.84) (63.46) (46.47)
-No Rule 25.85 35.06 74.78 74.14 72.38
+No Rule 35.06 25.85 74.78 74.14 72.38
 (59.92) (63.56) (23.56) (6.84) (23.38)
-Random 52.31 55.34 83.47 79.62 73.83
+Random 55.34 52.31 83.47 79.62 73.83
 (99.13) (115.59) (49.19) (44.98) (50.51)
-Back 48.52 49.71 85.21 70.0 74.41
+Back 49.71 48.52 85.21 70.0 74.41
 (92.64) (103.25) (19.86) (36.22) (18.17)
-Front 54.98 56.95 89.56 83.45 78.19
+Front 56.95 54.98 89.56 83.45 78.19
 (91.04) (100.38) (33.55) (39.17) (35.12)
 Table 2: Performance across 5 different reasoning tasks, evaluated with different sentence position removal.
-Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
+Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
 In Section 3, we demonstrated that gradually removing sentences from random and back positions can degrade model performance.
 To further explore this finding and assess the robustness of removing sentences from the front position, we conducted experiments with unguided random sentence removal (denoted as No Rule) and applied VARR+ with random position and reverse sentence order (denoted as Random and Back, respectively).
-As shown in Table 2, unguided random sentence removal resulted in a 25.30% decrease in performance relative to our method, highlighting the critical role of systematic configuration even after selecting sentences as the units of reduction.
-Additionally, Random and Back strategies result in an average 7.50% performance degradation compared to our method.
-This finding again indicates that earlier sentences in the reasoning path tend to be more redundant, and prioritizing their removal effectively balances rationale reduction while maintaining reasoning performance.
-5.5.3 Reinitializing the Optimizer
-G8K MQA CQA TQA SQA
-w/o reinit 45.18 48.70 87.82 83.72 76.74
-(115.91) (92.86) (32.95) (38.27) (33.89)
-w/ reinit 54.98 56.95 89.56 83.45 78.19
-(100.38) (91.04) (33.55) (39.17) (35.12)
-Table 3: Performance Comparison before and after the reinitializing optimizer in every training epoch. Each row’s first line denotes accuracy, while the second line shows the average generated tokens.
-We reinitialized the optimizer after each training epoch to stabilize training, inspired by Deng et al.
-(2024b).
-We employed the AdamW Loshchilov and Hutter (2019) optimizer, where the first and second moments are gradually updated based on current gradients, which could lead to unstable training in our framework.
-Specifically, after applying VARR throughout the dataset, a reduction in the rationale for some data between epochs could lead to instability in the training process.
-With a 19.32% performance gain achieved by reinitializing the optimizer, we reconfirm that Deng et al.
-(2024b)’s findings are robust across diverse datasets with complex semantic and syntactic structures in reasoning paths.
-5.5.4 Varying the Warm-up Steps
-Figure 4: Accuracy and Generated Token difference across various warm-up stages.
+As shown in Table 2, unguided random sentence removal resulted in a 25.30% decrease in performance relative to our method, highlighting the critical role of the verbosity evaluation even after selecting sentences as units of reduction.
+Furthermore, Random and Back strategies exhibited an average 7.50% performance degradation relative to our method.
+These results further support our observation that earlier sentences in the reasoning path tend to contain more redundancy, and their prioritized removal effectively balances rationale reduction while maintaining reasoning performance.
+G8K SQA TQA
+Exp-CoT 55.84 74.70 82.94
+(122.54) (46.47) (63.46)
+RAND-SENT REMOVAL 25.85 72.38 74.14
+(63.56) (23.38) (6.84)
+FRONT-SENT REMOVAL (1) 38.36 72.96 79.74
+(87.43) (24.34) (22.73)
+FRONT-SENT REMOVAL (2) 23.50 73.83 76.79
+(47.64) (9.97) (8.32)
+VARR+ 54.98 78.19 83.45
+(100.38) (35.12) (39.17)
+Table 3: Performance across different front sentence removal strategies.
+RAND-SENT-REMOVAL refers to the Random strategy, while ()($\cdot$)( ⋅ ) in the FRONT-SENT REMOVAL indicates the number of enforcedly removed front sentences.
+Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
+To further validate whether VARR enhances robustness compared to enforced removal of front sentences, we conducted additional experiments where the first and second sentences were enforced to be removed during the initial two training epochs.
+As shown in Table 3, enforced removal of front sentences leads to substantial performance degradation, while VARR demonstrates its contribution to preserving robustness.
+5.5.3 Varying the Warm-up Ratio
+Figure 4: Accuracy (barplot) and the average generated token (marker) across various warm-up stages on TriviaQA and CommomsenseQA.
 We evaluated our method against various warm-up stages on the TriviaQA and CommonsenseQA datasets.
@@
 These results suggest that 0.1 training steps provide sufficient time to inject reasoning abilities while enabling the systematic reduction of redundant reasoning sentences during the learning process.
+The zero warm-up configuration demonstrates comparable performance, though 0.1 warm-up steps are slightly better (detailed in Appendix F).
 Therefore, we set 0.1 as the default setting for the warm-up steps.
 5.5.4 Removal Ratio Analysis
-Figure 5: Removal ratio of redundant sentences throughout the training stage.
-The Removal Ratio, indicated on the y-axis, is calculated by dividing the number of sentences actually removed by the maximum potential removal ratio (i.e., ${size(\{r_{i}\})/r(t)})$.
-The x-axis represents the index of the epochs.
+Figure 5: Removal ratio of redundant sentences during training.
+The y-axis shows the removal ratio, calculated as the number of removed sentences divided by the maximum potential removal sentences (${size(\{r_{i}\})/r(t)}$).
+The x-axis represents training epochs.
 In Figure 5, we analyze the actual amount of rationale sentences removed during training.
@@
 Notably, a significant proportion of redundant sentences are removed in the early stages of training, with fewer sentences being removed as the model progresses through the middle to later stages, thereby stabilizing its training.
-A similar trend is observed across other datasets, as detailed in Appendix E.
+A similar trend is observed across other datasets, as detailed in Appendix G.
 6 Conclusion
-In this work, we propose sentence-level rationale reduction framework VARR and provide empirical evidence that models trained with non-redundant rationales achieve greater efficiency.
-We address the lack of principled criteria for identifying redundant sentences during the training phase by developing a reduction method that not only preserves the model’s reasoning capabilities but also reduces the likelihood of generating incorrect answers.
+In this work, we propose the sentence-level rationale reduction framework VARR and empirically demonstrate that models trained with non-redundant rationales achieve enhanced efficiency.
+We address the lack of principled criteria for identifying redundant sentences during training by developing a reduction framework that not only preserves the model’s reasoning capabilities but also reduces the likelihood of generating incorrect answers.
 Our experiments show that VARR can efficiently handle a diverse range of tasks with fewer generated tokens, without sacrificing its accuracy.
-Further empirical findings confirm that sentences are appropriate reasoning units for rationale reduction, highlighting the necessity of principled criteria for robust application.
-This work contributes novel insights to rationale reduction research, contributing to the efficient elicitation of reasoning in language models.
+This work contributes novel insights to rationale reduction research, contributing to the efficient reasoning elicitation in language models.
 Limitations
-While our work provides novel insights into rationale reduction research, our experiments were primarily conducted using a relatively small large language model, constrained by computational costs.
+While our work provides novel insights into rationale reduction research, our experiments were primarily conducted using a relatively small large language model and limited batch size, constrained by computational costs (i.e., a single A100-80GB GPU).
 Additionally, for the same reasons, it was not feasible to test the model with datasets featuring long sequences in both queries and rationales (Reddy et al., 2024; Yu et al., 2024a).
-However, given that the VARR/VARR+ frameworks are designed with a systematically structured approach, we believe that they could also be effective in scaled-up settings.
+Nevertheless, given the systematic design principles underlying the VARR/VARR+ frameworks, we believe their effectiveness would extend to larger-scale implementations.
+Furthermore, we reserve the application of VARR/VARR+ in iterative reasoning path generation and refinement/reflexion-based evaluation discussed in Section 2.1 for future work.
+Ethical Considerations
+Our work explores how LLMs can maintain their reasoning performance while improving efficiency.
+To this end, we conducted verbosity-aware rationale reduction (i.e., reasoning sentence pruning)-based CoT fine-tuning, requiring computational resources comparable to standard CoT fine-tuning.
+Additionally, we used only open-source LLMs and publicly available reasoning datasets with minimal preprocessing using gpt4o-mini’s (OpenAI, 2024) API.
+Therefore, we do not anticipate significant ethical issues arising from our work.
+On the contrary, we believe future works could leverage our analysis to reduce computational overhead in CoT inference settings.
+Acknowledgements
+We sincerely appreciate to Keonwoo Kim and Hyowon Cho for their sharp feedback during writing, Sangyeop Kim, Yukyung Lee, and Minjin Jeon for their continuous encouragement and inspiration during our study sessions, and Eunbi Choi for her invaluable help during the rebuttal process.
+This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.2018-0-00584, (SW starlab) Development of Decision Support System Software based on Next-Generation Machine Learning), National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2024-00335873), and National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2023-00217286).
 Appendix A Additional Datasets Analysis
+As shown in Section 3, the NLL tends to rise with increasing size of $\{r_{i}\}$ in random and back configurations, compared to the front configuration, as illustrated in Figure 6, this trend is consistent across both the GSM8K and StrategyQA datasets.
 Figure 6: NLL differences across varying sizes of $\{r_{i}\}$.
 The ‘original’ represents the NLL for the full rationale, while ‘front’, ‘random’, and ‘back’ indicate that ${r_{i}}$ are sampled from the front, random, and back indices of the full index set, respectively.
-As shown in Section 3, the Negative Log-Likelihood (NLL) tends to rise with increasing size of ${r_{i}}$ in random and back configurations, compared to the front configuration, as illustrated in Figure 7 this trend is consistent across both the GSM8K and StrategyQA datasets.
 Appendix B VARR Implementation
@@
 34:end for
-Appendix C Experimental Details
+Appendix C Additional Experimental Details
 For all experiments, we employ the AdamW optimizer Loshchilov and Hutter (2019), configured with a weight decay of 0.005.
 For the Mistral 7B model, we utilize an effective batch size of 12 with gradient accumulation set to 3, while the smaller Llama3.2 models use an effective batch size of 15.
+For Coconut (Hao et al., 2024), we set the max_latent_stage to 5 while maintaining all other hyperparameters as their default repository values unless otherwise mentioned.
 A constant learning rate of $5\times 10^{-6}$ is applied across all datasets, with bfloat16 precision.
 For Multiple-Choice and True/False tasks, complete sets of non-correct labels are employed to configure in-batch negatives to enhance the stability of the verbosity evaluation.
-To ensure a fair comparison, all baselines and methods are executed on a single A100 GPU with 80GB of memory for up to 5 epochs or 36 hours, whichever is reached first.
+To ensure a fair comparison, all baselines and methods are trained on a single A100 GPU with 80GB of memory for up to 5 epochs or 36 hours, whichever is reached first and experimented with single run evaluation (note that our setting is different from prompting/inference only setting).
+Regarding licensing, Mistral 7B is licensed under Apache License, Version 2.0, while Llama3.2 is governed by the Llama 3.2 Community License.
 Appendix D Details and Statistics of Datasets
 Dataset Reasoning Task Source Answer Format # TRAIN. # VALID. # TEST.
-GSM8K (Cobbe et al., 2021a) Arithmetic Cobbe et al. (2021b) Number 6999 472 1319
+GSM8K (Cobbe et al., 2021a) Arithmetic Cobbe et al. (2021b) Number 7000 473 1319
 MathQA (Amini et al., 2019a) Arithmetic Amini et al. (2019b) Multi-choice 29837 4475 2985
@@
 CommonsenseQA (Talmor et al., 2019) Multi-choice Kim et al. (2023) Multi-choice 609 38 115
-StrategyQA (Geva et al., 2021) Commonsense Sileo (2024) T/F 832 114 344
+StrategyQA (Geva et al., 2021) Commonsense Sileo (2024) T/F 1832 114 344
 Table 4: Comprehensive statistics of the datasets used in our experiments are provided.
@@
 (2023) and Sileo (2024) to access complete reasoning paths.
+GSM8K, StrategyQA, and CommonsenseQA are licensed under the MIT License, whereas MathQA and TriviaQA are distributed under the Apache License, Version 2.0.
 For our experimental analysis, we carefully selected a diverse set of five datasets used in prior works (Deng et al., 2023, 2024a; Liu et al., 2024; Yu et al., 2024b; Yin et al., 2024).
-To ensure explicit sentence boundaries, all datasets were processed using gpt4o-mini (OpenAI, 2024) to establish clear sentence demarcation (e.g., ‘He bikes 20*2=<<20*2=40>>40 miles each day for work So he bikes 40*5=<<40*5=200>>200 miles for work’ becomes ‘He bikes 202=<<202=40>>40 miles each day for work.
+To ensure explicit sentence boundaries, all datasets were preprocessed using gpt-4o-mini (OpenAI, 2024) to establish clear sentence demarcation (e.g., ‘He bikes 20*2=<<20*2=40>>40 miles each day for work So he bikes 40*5=<<40*5=200>>200 miles for work’ becomes ‘He bikes 202=<<202=40>>40 miles each day for work.
 So he bikes 405=<<405=200>>200 miles for work’).
 Table 4 comprehensively outlines each dataset, including its source and the size of the training, validation, and test samples.
-Appendix E Removal Ratio Analysis on additional datasets.
+MQA G8K CQA TQA SQA
+w/o reinit 48.70 45.18 87.82 83.72 76.74
+(92.86) (115.91) (32.95) (38.27) (33.89)
+w/ reinit 56.95 54.98 89.56 83.45 78.19
+(91.04) (100.38) (33.55) (39.17) (35.12)
+Table 5: Performance comparison before and after the reinitializing optimizer in every training epoch.
+Each row presents accuracy in the first line, with average generated tokens shown in parentheses in the second line.
+Appendix E Reinitializing the Optimizer
+We reinitialized the optimizer after each training epoch to stabilize training, inspired by Deng et al.
+(2024b).
+Our implementation uses the AdamW optimizer Loshchilov and Hutter (2019), where the first and second moments are gradually updated based on current gradients.
+Consequently, when VARR reduces rationales for certain data points between epochs, the training process could become unstable.
+As shown in table 6, the average 19.32% performance improvement achieved through optimizer reinitialization extends the findings of Deng et al.
+(2024b) beyond simple tasks (e.g., multiplication; synthesized dataset) to demonstrate effectiveness across diverse datasets with complex semantic and syntactic reasoning structures.
+Appendix F Zero Warm-up Stage
+warm-up 0.0 0.1 0.2 0.3 0.4
+TQA 82.01 83.72 83.45 81.19 81.97
+(40.11) (38.27) (47.10) (46.55) (47.97)
+CQA 90.34 89.56 87.82 87.82 86.08
+(32.81) (35.07) (39.38) (40.98) (43.61)
+Table 6: Accuracy and average generated tokens across the various warm-up stages.
+To validate whether the warm-up stage sensitivity could limit applicability, we conducted additional experiments extending warm-up stages to zero.
+As shown in Table, the zero warm-up stage shows comparable effectiveness and efficiency to the 0.1 warm-up stage, demonstrating the robustness of our method across different application scenarios.
+These findings suggest a zero-warm stage can be a good option for various application scenarios, but starting with a 0.1 warm-up stage can boost its stability.
+Appendix G Removal Ratio Analysis on additional datasets.
 Figure 7: Additional analysis of the Removal ratio on other datasets.
@@
 The x-axis represents the index of the epochs.
-In the GSM8K, Strategy-QA, and Math-QA datasets, similar trends are observed as in Section 5.5.5.
-Notably, in Strategy-QA, the Removal Ratio converges more rapidly than in other datasets.
+In the GSM8K, StrategyQA, and MathQA datasets, similar trends are observed as in Section 5.5.4.
+Additionally, we observe that our method reduces fewer sentences in later training stages, demonstrating that verbosity preserves essential reasoning sentences even when larger $r(t)$ encourages substantial reduction.
+Notably, in StrategyQA, the Removal ratio converges more rapidly than in other datasets.
 This is attributed to the dataset’s characteristics, where the sentences within the rationale predominantly list simple information rather than forming complex, interrelated structures.
 Consequently, redundant sentences are rapidly removed, leading to a rapid decrease in further removal activities.
-Appendix F Qualitative Analysis
+Appendix H Exploration on Complex Task
+Exp-CoT VARR+
+Acc 9.01 10.27
+Gen Tok 300.17 252.00
+Table 7: Performance comparison between Exp-CoT and VARR+ on MATH datasets.
+To further validate that our method remains effective on complex tasks, we conducted additional experiments with the MATH (Hendrycks et al., 2021) dataset.
+Considering computational constraints, we selected 500 samples from the math subset of OpenMathInstruct-2 (Toshniwal et al., 2024) for the training set and evaluated on the official test set.
+As described in Table 8, our results on MATH further demonstrate the versatility and future applicability of VARR+ on more complex tasks.
+Appendix I Training Time Analysis
+Dataset GSM8K (7K) TQA (2K)
+Training Time ($\uparrow$) 0.18 0.30
+Token Efficiency ($\downarrow$) 0.81 0.61
+Accuracy ($\uparrow$) 0.98 1.01
+Table 8: Training time analysis on GSM8K and TQA. All scores are calculated as the value for Explicit-CoT divided by the value for VARR+ (i.e., Explicit-CoT/VARR+).
+As VARR computes the likelihood of each sentence to ensure robust performance, training time inevitably increases as dataset size grows.
+We therefore quantitatively measure its efficiency on GSM8K and TQA relative to Explicit-CoT.
+Although there is a trade-off in training time, VARR+ improves token efficiency while preserving accuracy.
+Since training occurs only once while the improved efficiency benefits persist continuously afterward in test-time, we believe the additional training cost is justified.
+Appendix J Qualitative Analysis
 Table 9 illustrates the generation outputs for each method on the GSM8K test set.
-Explicit-CoT produces a complete reasoning path for answer generation; however, its redundancy in generating ‘needs a 3 piece place setting’ leads to hallucinations within the sentence, resulting in the incorrect final answer generation.
-In contrast, ICoT-SI, which removes reasoning paths at the token-level, results in a generation that lacks sufficient reasoning capabilities.
-Conversely, VARR+ efficiently generates key information by producing ‘\$15.00’ in a single sentence, leading to more effective and efficient generation.
+While Explicit-CoT generates a complete reasoning path, its redundant generation of ‘needs a 3 piece place setting’ introduces hallucinations into the reasoning process, ultimately leading to an incorrect answer.
+ICoT-SI’s token-level elimination of reasoning paths, on the other hand, results in generations that lack sufficient reasoning capacity.
+Additionally, Coconut’s predefined hidden-states fail to compress the complete reasoning path in its token-level compression approach.
+In contrast, VARR+ demonstrates efficient generation by concisely producing ‘\$15.00’ in a single sentence, achieving both effectiveness and efficiency in its output.
+Additional qualitative examples on the MathQA dataset are outlined in Table 10.
 Question
@@
 All total, the place settings will cost \$216 for dinner plates, \$180 for bowls and \$144 for salad plates for a total of 216+180+144 = \$540.00
-### 540
+#### 540
 ICoT-SI
 each so 12*6 = \$72.00 She’s buying 3 pieces of place settings (dinner plates, salad plates and bowls) for \$72.00 each so 3*72 = \$216.00
-### 216
+#### 216
+Coconut
+<|start-latent|><|latent|><|latent|><|latent|><|end-latent|>
+#### 4.00
 VARR+
 She needs 12 place settings and each setting costs \$15.00 so she will spend 12*15 = \$180.00
-### 180
+#### 180
 Ground Truth Answer
-### 180
+#### 180
 Table 9: Qualitative Example: Generation outputs on GSM8K test set.
+Question
+The cost price of 20 articles is the same as the selling price of $x$ articles. if the profit is 25%, find out the value of x.
+a) 13, b) 14, c) 15, d) 16, e) 17
+Explicit-CoT
+explanation : let the cost price 1 article = rs . 1
+cost price of $x$ articles = rs . $x$
+selling price of x articles = 125% of $x$ = rs . ( 125 / 100 $x$ ) = rs . ( 25 / 20 $x$ ) = rs . 5 / 4 $x$
+= > $x$ = 20
+cost price of 20 articles = rs . 20 selling price of 20 articles = rs . ( 25 / 4 $\times$ 20 ) = rs . ( 25 / 4 $\times$ 20 ) / 20 = rs . 1
+= > $x$ = 20 answer : option a
+#### a
+ICoT-SI
+of $x$ articles = $x$
+s.p of $x$ articles = rs . 20
+gain = rs . ( 20 - $x$ )
+gain % = 25 % of $x$ = rs . ( 20 - $x$ ) / $x$$\times$ 100 = 25
+= > 20 - $x$ = 5 $x$ = > 4 $x$ = 20 - $x$ = > 5 $x$ = 20 - $x$ = > 20 = 5 $x$ = > 200 = 5 $x^{2}$ = > 400 = 5 $x^{2}$ = > 400 / 5 = $x^{2}$ = > $x^{2}$ = 80 = > x = 10 or 10 . c
+#### c
+Coconut
+<|start-latent|><|latent|><|latent|><|latent|><|end-latent|>
+#### c
+VARR+
+explanation : let the cost price of $x$ articles = $x$
+s.p of $x$ articles = 20
+20 = 1.25 $x$ = > $x$ = 16 option d.
+#### d
+Ground Truth Answer
+#### d
+Table 10: Qualitative Example: Generation outputs on MathQA test set.